### v2

# Unlock the Power of Frozen LLMs in Knowledge Graph Completion

Abstract
Traditional knowledge graph completion (KGC) methods rely solely on structural information, struggling with the inherent sparsity of knowledge graphs (KGs). Large Language Models (LLMs) learn extensive knowledge from large corpora with powerful context modeling, making them promising for mitigating the limitations of previous methods. Directly fine-tuning LLMs offers great capability but comes at the cost of huge time and memory consumption, while utilizing frozen LLMs yields suboptimal results. In this work, we aim to leverage LLMs for KGC effectively and efficiently. We capture the context-aware hidden states of knowledge triples by employing prompts to stimulate the intermediate layers of LLMs. We then train a data-efficient classifier on these hidden states to harness the inherent capabilities of frozen LLMs in KGC. Additionally, to reduce ambiguity and enrich knowledge representation, we generate detailed entity descriptions through subgraph sampling on KGs. Extensive experiments on standard benchmarks demonstrate the efficiency and effectiveness of our approach. We outperform traditional KGC methods across most datasets and, notably, achieve classification performance comparable to fine-tuned LLMs while enhancing GPU memory efficiency by 188×188$\times$188 × and accelerating training and inference by 13.48×13.48$\times$13.48 ×.
Unlock the Power of Frozen LLMs in Knowledge Graph Completion
Bo Xue, Yi Xu, Yunchong Song, Yiming Pang, Yuyang Ren, Jiaxin Dingthanks:  Jiaxin Ding is the corresponding authors., Luoyi Fu, Xinbing Wang Shanghai Jiao Tong University, Shanghai, China {sappho_x, jiaxinding, yiluofu} @sjtu.edu.cn
1 Introduction
Knowledge graph completion has become a crucial endeavor to improve the application of knowledge graphs (Chen et al., 2020), which aims to complete the knowledge graph and expand its scale by predicting the potential relationship between existing entities and discovering new relational facts (Shen et al., 2022b). Traditional knowledge graph embedding methods only utilize structure information from the knowledge graph (Bordes et al., 2013; Trouillon et al., 2016; Sun et al., 2019), which suffer from the sparseness of KGs (Yao et al., 2019). Leveraging the powerful contextual modeling capabilities and extensive stored world knowledge of LLMs can mitigate this challenge by utilizing the rich semantics within knowledge graphs, thereby enhancing the performance of knowledge graph completion (Wei et al., 2023).
These approaches generally fall into two types: using frozen LLMs with prompting techniques and fine-tuned LLMs specifically for KGC. The first type leverages frozen LLMs with prompt techniques, such as in-context learning. These techniques enable models to perform KGC directly through natural language outputs (Wei et al., 2023; Yao et al., 2023), harnessing the pre-trained knowledge embedded within these models without additional training. While this approach offers computational efficiency, it may not fully capture the task-specific nuances required for optimal KGC performance. The second type involves fine-tuning the LLMs using supervised techniques, including instruction tuning, to adapt models specifically for KGC tasks. This method involves fine-tuning LLMs, such as LLaMA, through the use of training samples (Yao et al., 2023), which can lead to improved performance by aligning the models more closely with the KGC objectives (Zhang et al., 2023; Yao et al., 2023). However, fine-tuning demands significant increases in time and resource expenditure. This dichotomy raises a fundamental question: How can we utilize LLMs to accomplish KGC tasks with both efficacy and efficiency?
To answer this question, we explore the reasons behind the underperformance of frozen LLMs compared to their fine-tuned counterparts. First, while LLMs contain vast amounts of world knowledge encoded within their parameters, accessing the knowledge and decoding it into responses effectively through frozen models can be challenging (Wang et al., 2023). LLMs can produce hallucinations-plausible but incorrect outputs (Zhao et al., 2023)-making them unreliable for precise KGC tasks. Fine-tuning directly modifies the models’ weights through task-specific data enhancing the retrieval of relevant knowledge and improving performance (Zhou et al., 2024). Second, KGs often involve complex relationships and ambiguous entity representations (Li et al., 2024), which can be difficult for a frozen LLM to handle such nuances accurately without specific tuning. Fine-tuned models can be explicitly trained to interpret and disambiguate the complexities enabling better linkage of entities within KGs.
To overcome the limitations of frozen LLMs in KGC tasks, we introduce probing techniques that utilize stimulation prompts, complemented by detailed entity descriptions, as shown in Figure 1. These methods are designed to effectively extract and leverage the latent knowledge embedded within the models’ parameters, and minimize ambiguous entity representations, thereby enhancing its application in KGC without the need for extensive retraining. Specifically, our research concentrates on the core tasks of knowledge graph completion: triple classification and relation prediction.
We stimulate the intermediate layers of LLMs to capture the context-aware hidden states of knowledge triples. Moreover, we improve the precision of knowledge representation by aligning KG entities with the LLMs, reducing the ambiguity of knowledge triples through the generation of entity descriptions via subgraph sampling from KGs. We train a data-efficient classifier on the obtained hidden states suppervised by labels, to fully harness the inherent KGC capabilities of the frozen LLMs.
Figure 1: The overall architecture of our proposed probing method for the triple classification. Constructing positive and negative sample pairs, followed by stimulating the middle layers of the language model to obtain hidden states for completing the KGC task. Entity descriptions are generated by the subgraph entity description generator. The only component that needs training in the entire architecture is the data-efficient classifier used to classify hidden states, which can be MLP, logistic regression, etc.
Our contributions can be summarized as follows.
  • We introduce a novel probing method tailored for KGC tasks, designed to stimulate language models to predict knowledge graph triples directly. We further generate knowledge graph entity descriptions through the use of subgraphs, which are then seamlessly integrated into our probing techniques.
  • Our approach employs LLMs with both effectiveness and efficiency to achieve KGC. By utilizing only a subset of the dataset, our method achieves a relative performance improvement of up to 2.7% compared to fine-tuning the LLM on the entire training set. Additionally, it significantly reduces GPU memory usage by a factor of 188 during the training phase and decreases time consumption by a factor of 13.48, all while maintaining comparable performance. [^1] For a detailed comparison of efficiency, please refer to Table 8 in the Appendix B.3.
  • We conduct extensive experiments on six standard knowledge graph datasets to benchmark our method against multiple baseline approaches. We investigate the versatility of our method with various language models, assessing its efficacy and data efficiency through rigorous experimental validation.
2 Method
Our method is designed based on an analysis of the two primary reasons for the underperformance of frozen LLMs compared to fine-tuned counterparts. First, frozen LLMs may not effectively access the necessary knowledge to generate accurate responses. We demonstrate in Section 2.2 how to achieve this for the frozen LLM using probing techniques. Second, KGs often involve ambiguous entity representations, which can be difficult for LLM to handle accurately. In Section 2.3 we show how to achieve this with the generation of entity descriptions.
2.1 Formulation and Notations
A knowledge graph, $G$, comprises a set of entities, $E$, and relations, $R$. It can be represented as a set of triples $G=\{(h,r,t)\}$, where $h\in E$ is the head entity, $t\in E$ is the tail entity, we collectively refer to the head entity and tail entity as $e$, and $r\in R$ represents the relation between them. $D$ is the set of descriptions for each entity and relation. We denote $D(e)$, $D(r)$ as the textual description of each entity and each relation.
We focus on triple classification and relation prediction. Triple classification involves evaluating the plausibility of a given knowledge triple, to determine whether the triple is true or not and relation prediction predicts the missing relation in a triple.
2.2 Stimulating the Intermediate Layers of Large Language Models
Knowledge graphs consist of numerous triples, and the format of these triples is challenging for the LLM to understand, which may mislead the LLM to retrieve irrelevant knowledge and reveal the wrong responses. To enhance the performance of LLMs in KGC, we leverage the internal judgment mechanisms of the language model (Zou et al., 2023). Our method involves strategically triggering the necessary depth of knowledge embedded within the LLM to elicit the correct responses required for accurate KGC, thereby improving the overall performance of the KGC task.
Specifically, since different layers of the language model have different roles (Azaria and Mitchell, 2023), we determine the most suitable layer for the task by experimenting with the validation set. For the selection of intermediate layers, see Appendix B.5. First, we construct prompt templates to stimulate the model to generate the required classification information in the intermediate layers. In the triple classification, our goal is to determine whether the triples in the knowledge graph are true. Therefore, we use the following prompt template, $PT_{1}$, to create positive and negative sample pairs for training, as shown in Figure 2. Notice that the prompt does not describe positive or negative information for triples.
Then, we use existing triples in the knowledge graph to construct positive samples $S^{+}$, and employ the negative sampling technique (Lv et al., 2022) to obtain negative samples $S^{-}$. Finally, we obtain $N$ training data samples. For each training text, we input it into the language model to obtain the representation of each text in the intermediate layers. We can obtain information about the plausibility of triples from the hidden state corresponding to the token immediately preceding the model prediction, which is the last token of the input text (Zou et al., 2023). Specifically, we capture plausibility information for positive samples $s^{+}\in S^{+}$ by:
$v\left(s^{+}\right)=\{M_{1:i}\left(s^{+}\right)[-1],0<i<L\},$ (1)
where $M$ represents the language model, $M_{1:i}$ represents the first $i$ layers of the model, $[-1]$ refers to the hidden state corresponding to the last token, $L$ denotes the number of layers in the model. The dimension of $M_{1:i}(s^{+})[-1]$ is the embedding dimension of the language model. Negative samples follow the same principle.
At last, we train a classification model using $v(s^{+})$ and $v(s^{-})$ to distinguish between plausible and implausible triples. The goal is to learn a boundary that can separate the hidden states of the intermediate layers of the LLM corresponding to positive samples (which are true) from those corresponding to negative samples (which are false). We can use classification models such as multilayer perceptron (MLP), SVM, etc., for classification training. Taking logistic regression to determine the plausibility of samples as an example:
$P\left(Y=\mathbf{1}|x\right)=\frac{1}{1+e^{-\left(w^{T}x+b\right)}},$ (2)
where x represents $v(s)[l]$, $l$ is the layer index, $w$ and $b$ are weights and bias to be learned.
For relation prediction, we transform the binary classification into multi-class classification. By utilizing the prompt in Figure 7, we stimulate the LLM to generate hidden states that capture the characteristics of different relation types in the KG. Details can be found in the Appendix B.1.
By focusing on extracting classification results directly from the model’s internal processing, we ensure that the responses are both relevant and precise, thereby improving the effectiveness of the KGC task.
Figure 2: Prompt templates to stimulate the LLM for triple classification.
2.3 Subgraph Entity Description Generator
The semantics of entities in the knowledge graph are complex, making it challenging for the language model to interpret based solely on the entity’s name due to inherent ambiguity and polysemy. Therefore, we inform the language model about the meaning of each entity in the input. The meaning of each entity in KG is generated by its one-hop subgraph.
Specifically, we extend the prompt template constructed in Section 2.2, aiming to assist the language model in aligning internal entities with entities in the knowledge graph. For triple classification, since we employ the probing method to extract KGC features from the samples, we need to ensure that the language model can be stimulated to produce effective hidden states that can be used by the classifier to predict the plausibility of triples. Therefore, we construct a new prompt $PT_{2}$, as shown in Figure 2. By introducing the $D(e)$, LLMs are able to generate more precise and distinguishable hidden states, thereby improving the overall classification performance.
To address incomplete entity descriptions within certain datasets, we have devised a method to generate $D(e)$. We construct $D(e)$ in two ways. The first approach involves a straightforward method where we concatenate all the triples within the one-hop subgraph surrounding the entity $e$, thereby forming $D(e)$ (Zhang et al., 2023). However, this direct concatenation can sometimes lead to ineffective or misleading stimuli for the LLM, as the structure of these subgraph triples may significantly differ from the text seen during the LLM’s pre-training phase. This divergence can adversely affect the model’s response accuracy. To mitigate this issue, our second approach leverages a language model to transform these triples into a format that is more comprehensible to the LLM. By rephrasing the triples from the one-hop subgraph around the entity $e$ into natural language text, we produce a more refined and model-friendly version of $D(e)$:
$D(e)=\operatorname{concat}\{\operatorname{transform}(\{h_{1}\}\{r_{1}\}\{t_{1}\}),$ (3)
$(h_{1},r_{1},t_{1})\in\operatorname{subgraph}(e)\},$
where $\operatorname{subgraph(\cdot)}$ is a function that retrieves the set of all triples in the one-hop subgraph where the entity resides; $\operatorname{transform(\cdot)}$ is a function that converts the triple into sentence; $\operatorname{concat(\cdot)}$ is a function to process triple statements. For the first approach, $\operatorname{concat(\cdot)}$ is a string concatenation function that joins each transformed triple statement with a separator. For the second approach, we need to utilize LLM with the in-context learning technique (Brown et al., 2020) to generate descriptions of entities based on the triples in their subgraph. Please refer to Appendix A.1 for details.
By generating long entity descriptions using the subgraph description generator, we align the internal understanding of entities in the LLM with the entity meanings in the KG, which is similar to the entity linking process in knowledge graphs. In this way, the LLM will better understand the semantics of the entities, and improve the KGC performance.
3 Experiments
In the experiments, our objective is to answer the following research questions (RQs):
  • RQ1: Can frozen LLMs with the proposed probing methods effectively achieve performance on par with their fine-tuned counterparts?
  • RQ2: How does our method perform across different language models and classification models?
  • RQ3: What are the data efficiency and computation efficiency of our proposed method for KGC?
3.1 Datasets
Dataset |E| |R| # Train # Valid # Test
FB13 75,043 13 316,232 5,908 23,733
WN11 38,696 11 112,581 2,609 10,544
FB15K-237N 13,104 93 87,282 14,082 16,452
WN18RR 40,943 11 86,835 6,068 6,268
UMLS 135 46 5,216 1,304 1,322
YAGO3-10 123,182 37 1,079,040 5,000 5,000
Table 1: Statistical information of datasets.
In our experiments, we employ five widely-used KG datasets for triple classification: FB13 (Socher et al., 2013), WN11 (Socher et al., 2013), FB15K-237N (Lv et al., 2022), WN18RR (Dettmers et al., 2018), and UMLS (Yao et al., 2019), as well as a commonly used KG dataset for relation prediction: YAGO3-10 (Dettmers et al., 2018).
For FB13, we use Wikipedia entity descriptions as input sentences. For FB15K-237N, the entity descriptions provided by (Xie et al., 2016) are used. Synset definitions from (Yao et al., 2019) served as entity descriptions for WN18RR. In the case of UMLS, entity descriptions provided by the dataset are used. In the following, we refer to the entity descriptions as non-generated entity descriptions.
3.2 Experimental Settings
To generate entity descriptions, for the first approach in Section 2.3, we concatenate all subgraph triples to generate descriptions. For the second approach, we employ GPT-3.5-turbo as the LLM. Due to cost considerations, we randomly select 4000 test examples from each dataset to generate entity descriptions. We mark the results of these methods with an asterisk (*). In the following text, we refer to these entity descriptions as generated entity descriptions (Tri)/(GPT).
We compare our method with four structure information-based methods: TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), ComplEx (Trouillon et al., 2016) and RotatE (Sun et al., 2019), as well as seven additional information-based methods: KG-LLAMA (Yao et al., 2023), KG-BERT (Yao et al., 2019), KG-T5 (Saxena et al., 2022), LLaMA-7b, Alpaca-7B, LLaMA-7B-ICL and Alpaca-7B-ICL (Zhang et al., 2023). KG-LLAMA, KG-BERT and KGT5 fine-tune the language models using all the training samples. LLaMA-7b and Alpaca-7B using the frozen language model to predict. LLaMA-7B-ICL and Alpaca-7B-ICL using 2-shot samples with in-context learning to predict. Please refer to Section 4.1 for detailed introductions of these methods.
For structure information-based methods, we reproduce the results with OpenKE (Han et al., 2018). For additional information-based methods, we reproduce the results with official code implementation. For KGT5, there’s no public record or code for the method on the FB15K237N, WN18RR, UMLS dataset. Please refer to Appendix A.2 for detailed experimental settings.
Method Pred Samples FB13 WN11 15K237N WN18RR UMLS
TransE S-F ALL 0.815 0.759 0.697 0.884 0.845
DistMult S-F ALL 0.862 0.871 0.587 0.851 0.864
ComplEx S-F ALL 0.857 0.862 0.657 0.841 0.908
RotatE S-F ALL 0.819 0.847 0.685 0.882 0.921
KG-BERT L-MLP ALL 0.904 0.935 0.560 0.927 0.773
KGT5 N-L ALL 0.663 0.728 - - -
LLAMA-7B N-L 0 0.699 0.661 0.573 0.458 0.658
LLAMA-7B-ICL N-L 2 0.501 0.500 0.578 0.502 0.538
KG-LLAMA-7B N-L ALL 0.892 0.955 0.748 0.921 0.858
KG-LLAMA-SAMPLED N-L 3k/10k 0.727 0.682 0.614 0.574 0.834
LLAMA-MLP H-MLP 3k/10k 0.855 0.872 0.674 0.858 0.882
LLAMA-DES-H H-MLP 3k/10k 0.901 - 0.738 0.931 0.862
LLAMA-DES2 (Tri) H-MLP 3k/10k 0.842 0.755 0.671 0.779 0.782
LLAMA-DES2 (GPT)* H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
Table 2: The main experiment results (accuracy) of triple classification. In the $Pred$ column, $S\mbox{-}F$ uses score function, $L\mbox{-}MLP$ uses MLP on last layer hidden states, $N\mbox{-}L$ uses natural language, and $H\mbox{-}MLP$ uses MLP on intermediate layer hidden states to predict. $Samples$ indicates training sample counts, with $3k/10k$ indicating $10k$ samples for the UMLS dataset and $3k$ for others. LLAMA-MLP uses LLaMA-7B with MLP for hidden state classification and $PT_{1}$ for stimuli generation; LLAMA-DES-H combines non-generated descriptions with $PT_{2}$ for stimulus generation; LLAMA-DES2 (Tri)/(GPT) uses first/second generated descriptions with $PT_{2}$ for stimuli generation. Top three results per dataset are bolded.
3.3 Main Results (For RQ1)
From Table 2, it can be observed that LLAMA-DES-H and LLAMA-DES2 (GPT), when using a partial training set, exhibit comparable performance to KG-LLAMA-7B, the latter being fine-tuned with all training triples. Moreover, when trained on subsets of equivalent size, LLAMA-DES-H and LLAMA-DES2 (GPT) markedly surpass KG-LLAMA-SAMPLED, the existing state-of-the-art approach that relies exclusively on LLaMA for KGC. For more details, including the results of baseline models such as Alpaca-7B-ICL and Structural-aware-IT, and the performance of our method on the full training set, refer to Appendix B.4.
The experimental results presented in Table 2 also validate the effectiveness of the approaches proposed in Section 2.2 and Section 2.3. Specifically, LLAMA-DES-H and LLAMA-DES2 (GPT) outperform LLAMA-MLP, demonstrating the effectiveness of using subgraph entity description generator to improve the LLM’s understanding of entities in the KG, as proposed in Section 2.3. Furthermore, LLAMA-MLP demonstrates superior performance compared to LLAMA-7B and LLAMA-7B-ICL, affirming the effectiveness of stimulation method proposed in Section 2.2. Thus, the experimental results substantiate the effectiveness of our proposed methodology. Additionally, the frozen LLM can also achieve similar KGC effects as that of fine-tuning the LLM.
Notice that the performance on the UMLS dataset is special, where the use of generated descriptions hinders the model’s classification accuracy, and the use of non-generated entity descriptions is also worse than not using any entity descriptions. It is postulated that the underlying reason may lie in the highly specialized and domain-specific nature of entity names in the UMLS dataset, which are often precise medical terms with limited room for multiple interpretations. The entities in the knowledge graph have well-aligned with the internal entities of the LLM. Meanwhile, using entity descriptions may be affected by noisy description text, and descriptions generated by GPT-3.5-turbo may be influenced by hallucination, thereby affecting the classification effectiveness of the induced hidden states.
In addition to the triple classification, we also validate the effectiveness of the proposed probing method on the relation prediction task as shown in Table 3. The experimental results similarly demonstrate that the probing method can achieve the nearly same performance as full training set fine-tuning, even when using only 0.6% of the training data. Detailed methodology and analysis are provided in Appendix B.1.
Figure 3: Differences in KGC performance across different layers of the LLAMA-DES2 (GPT) [left] and Mistral-DES2 (GPT) [right]. The horizontal axis represents the position of the intermediate layers in LLaMA or Mistral (which have different backbones), and the vertical axis represents the dataset.
Method Pred Samples YAGO3-10
KG-BERT L-MLP ALL 0.6816
KGT5 N-L ALL 0.5714
ChatGLM-6B N-L 0 0.0658
KG-ChatGLM-6B N-L ALL 0.5662
LLaMA-7B N-L 0 0.0348
LLaMA-13B N-L 0 0.0040
KG-LLaMA-7B N-L 0 0.7028
LLAMA-MLP-7B H-MLP 6996 0.5968
LLAMA-DES2 (GPT)* H-MLP 6996 0.6824
Table 3: The experiment results (Hits@1) of relation prediction. LLAMA-MLP-7B refers to using LLaMA-7B as the base language model, using MLP as the hidden state classifier, and generating stimuli using $PT_{3}$; LLAMA-DES2 (GPT) refers to using the second generated description combined with $PT_{4}$ to generate stimuli.
3.4 Ablation Study (For RQ2)
Through the analysis depicted in the left portion of Figure 3, it is observed that with an increase in the index of the intermediate layer, the efficacy of hidden states for triple classification demonstrates an initial incline followed by a subsequent decline. Intermediate layers closer to natural language text output generally exhibit better prediction performance compared to the lowest layers. This observation corroborates the notion posited by Geva et al. (2020) that language models encode knowledge differentially across their Feed-Forward layers, with higher layers integrating information from lower layers. During pretraining, the lower layers may not have stored certain knowledge, and thus are unable to generate hidden states conducive to accurate triple classification. Furthermore, the prediction performance of the layers closer to the natural language text output is generally worse than those in the middle layers. This phenomenon may be attributed to the hallucination issue inherent in LLaMA, wherein higher layers could potentially either omit or alter triple information, leading to a blurred distinction between positive and negative triple hidden states induced by stimuli (Jiang et al., 2024; Zou et al., 2023). Detailed explanations can be found in Appendix B.5. Akin observations are discernible on the right side of Figure 3 for the Mistral model, which has a different backbone from LLaMA. Meanwhile, as observed in Table 5, our method demonstrates strong versatility across different backbone models, achieving performance comparable to full dataset fine-tuning with only a small sample size.
Regarding classification models, a comparative analysis encompassing MLP, SVM, and Logistic Regression, as delineated in Table 4, is conducted. The result shows that MLP outperforms all three datasets compared to the other two classification models, while Logistic Regression also exhibits results close to MLP. This underscores the efficacy of our stimulus methodology, whereby hidden states of positive and negative samples within a high-dimensional space are easily distinguishable.
Method FB13 WN11 FB15K-237N
LLAMA-SVM 0.821 0.866 0.594
LLAMA-LR 0.839 0.860 0.662
LLAMA-MLP 0.855 0.872 0.674
Table 4: The variation in KGC prediction performance across different datasets using SVM, Logistic Regression, MLP as classification models. All these methods use LLaMA-7B as the base model, and generate stimuli using $PT_{1}$.
Model Method FB13 WN11
LLaMA KG-LLAMA-7B 0.892 0.955
LLAMA-MLP 0.855 0.872
LLAMA-DES-H 0.901 -
LLAMA-DES2 (GPT)* 0.912 0.917
Mistral KG-MISTRAL-7B 0.891 0.962
MISTRAL-MLP 0.846 0.901
MISTRAL-DES-H 0.880 -
MISTRAL-DES2 (GPT)* 0.893 0.925
Gemma KG-GEMMA-7B 0.897 0.944
GEMMA-MLP 0.806 0.893
GEMMA-DES-H 0.891 -
GEMMA-DES2 (GPT)* 0.867 0.917
Table 5: The variation in KGC prediction performance across different datasets using LLaMA, Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024) as base models (which have different backbones). Please refer to Table 2 for the number of training samples used by different methods.
3.5 Efficiency Study (For RQ3)
As illustrated in Figure 4, our method notably surpasses both KG-LLAMA and KG-BERT on the FB13 when trained with only 100 samples, as opposed to the 3000 samples used by the latter methods. For the UMLS, the performance of KG-LLAMA and KG-BERT with 3000 training samples is comparable to that of LLAMA-DES-H trained with only 500 samples. Additionally, combining the results from Table 2 and Table 10, for the FB13 dataset, our method attains 97.2% of the performance achievable with the full training set, utilizing only 500 samples, which constitutes a mere 0.079% of the total training samples. In contrast, KG-LLAMA and KG-BERT achieve only 78.7% and 63.1% of the maximum performance, respectively. For the UMLS, with 9.58% of the training set, our method reaches 90.6% of the full training set performance. For other datasets, please refer to Appendix B.6. Our experimental results confirm that the language model is capable of generating effective hidden states through the techniques described in Sections 2.2 and 2.3, demonstrating its potential for data-efficient scenarios within the context of KGC tasks.
Figure 4: The prediction performance of KG-LLAMA, KG-BERT, and LLAMA-DES-H on triple classification varies with the size of the training dataset. The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples.
Meanwhile, we also evaluate the compute efficiency of our method. Table 8 shows that our method can achieve significantly lower consumption compared to parameter-efficient fine-tuning even in the setting of the full dataset. Compared to fine-tuned LLMs, we boost the GPU memory efficiency to 188×188$\times$188 × during model training and speed up 13.48×13.48$\times$13.48 × overall. Please refer to Appendix B.3 for details.
4 Related Work
4.1 Knowledge Graph Completion
Knowledge graphs are typically characterized by their inherent incompleteness, often containing substantial amounts of implicit or missing valuable knowledge (Hogan et al., 2021). The prevailing methodologies in KGC can be categorized into two main classes: structure information-based methods and additional information-based methods (Shen et al., 2022b). Structure information-based methods leverage the inherent structural data present within knowledge graphs. For example, RESCAL (Nickel et al., 2011), DistMult (Yang et al., 2014), COMPGCN (Vashishth et al., 2019), and ConvE (Dettmers et al., 2018) primarily use scoring functions to evaluate the semantic similarities of entity or relation embeddings within a latent embedding space to infer missing information. Additional information-based methods incorporate a variety of supplementary data to enrich the knowledge graphs. These methods utilize node attributes, entity-related information, or relational path data to enhance the features of knowledge embeddings critical for KGC tasks. The prominent examples are KG-BERT (Yao et al., 2019) and LASS (Shen et al., 2022a), which represent entities, relations, and triples as textual sequences, turning KGC into a sequence classification problem. KG-S2S (Chen et al., 2022) and KGT5 (Saxena et al., 2022) employ encoder-decoder architecture language models to complete KGC in a generative manner. KG-LLAMA (Yao et al., 2023) utilizes triples composed of entity and relation names as sentences, parameter-efficient fine-tuning the LLaMA model for KGC. Li et al. (2024) enhances KGC with LLM by generating entity descriptions and designing loss functions for fine-tuning.
4.2 Large Language Models
Several studies have explored the mechanisms through which Large Language Models process, interpret, and generate factual content. As highlighted by Saier et al. (2023), these models function as extensive repositories of knowledge, encapsulating a vast array of information about the world within their parameters. This insight into LLMs underscores their potential as powerful tools for a wide range of applications, including knowledge graph completion, by leveraging the rich informational content stored in their intricate network structures. Ramrakhiyani et al. (2023) have conducted evaluations across multiple Large Language Models to assess their knowledge of world geography using probing techniques. Similarly, Onoe et al. (2022) developed a methodology for automatically constructing cloze prompts by masking out spans, aiming to assess the language models’ understanding of entities through factual probing. In addition, Azaria and Mitchell (2023) and Zou et al. (2023) have explored training classifiers to estimate the truthfulness of statements based on the activations of the hidden layers in LLMs.
5 Conclusion
In this paper, we introduce a novel data-efficient probing method that completes KGC by leveraging the internal knowledge of LLMs. Experiments show that using the probing method with generated entity descriptions, frozen LLMs can achieve performance comparable to that of their fine-tuned counterparts. Simultaneously, this method saves a significant amount of GPU memory and time compared to parameter-efficient fine-tuning approaches. This work explores the potential of frozen LLMs in KGC tasks with both efficacy and efficiency. Compared to fine-tuning LLMs, we hit the core of KGC problem in a better way.
Limitations
Our current architectural design is only applicable to open-source LLMs or closed-source models that provide access to token embeddings; it is not suitable for closed-source models that exclusively output text. Recognizing these limitations, future work will aim to expand our architecture to accommodate closed-source LLMs that only output text or token probabilities. Additionally, future efforts could integrate structure information-based KG models, such as TransE, with our methodology in a hybrid approach.
Appendix A Experimental Details
A.1 Generate Entity Descriptions using GPT-3.5-turbo
In Section 2.3, we utilize LLM with the in-context learning technique to generate descriptions for entities based on the triples in their one-hop subgraph. In the experiments, we utilize the GPT-3.5 as the LLM. The prompt template used to generate entity descriptions is shown in Figure 5. Specifically, the $EN(e)$ is the entity name of entity $e$; the $\operatorname{concat\_subgraph()}$ is a function that first converts triples from the one-hop subgraph where the entity is located into sentences, and then concatenates these sentences using a specified delimiter. It is important to note that all the triples used in the prompt are from the training set, preventing potential data leakage from the test set in the generated entity descriptions. Table 6 shows the descriptions of entities generated using this method, thus achieving "entity alignment" between the knowledge graph and the LLM.
Entity Generated Description Non-Generated Description
Gustav Mahler
Gustav Mahler was a Jewish composer
known for his significant contributions
to the world of music. He lived in the
United States and was of German nationality.
Gustav Mahler (German; 7 July 1860 – 18 May 1911)
was an Austro-Bohemian late-Romantic composer,
and one of the leading conductors of his generation…
James Hayter
James Hayter was a male actor born in Scotland,
known for his profession in acting.
He passed away in Spain.
Henry James Goodenough Hayter
(23 April 1907 – 27 March 1983), better known
as James Hayter, was a British actor. He is best
remembered for his roles as Friar Tuck in the film The Story
of Robin Hood and His Merrie Men (1952) and…
Table 6: Comparison of generated entity descriptions with non-generated entity descriptions on the FB13 dataset.
Figure 5: The prompt template used to generate entity descriptions based on the one-hop subgraph where the entity is located for GPT-3.5-turbo.
A.2 Implementation Details
Figure 6: The prompt template used in Alpaca-7B-ICL on the WN18RR dataset.
In our implementation, we experimented with these baselines using the hyperparameters recommended in their paper.
For the negative sampling employed on the datasets, we utilize the same negative sampling technique as Yao et al. (2019) for FB13 and WN11. Specifically, FB15K-237N is designed to extract hard negative triples, providing a more challenging evaluation and reducing false negatives in the validation/test datasets during dataset construction. A similar approach is adopted for constructing negative samples for WN18RR and UMLS. Meanwhile, for all experiments in our study, the training sets contain an equal number of positive and negative samples.
For the MLP classifier configuration, we opt for a batch size of 64, a learning rate of 3e-5, and utilize the AdamW optimizer, with training restricted to 30 epochs. For LLaMA-7B-ICL and Alpaca-7B-ICL, we randomly select one positive sample and one negative sample from the training set of each dataset, respectively. The prompt for KGC using Alpaca-7B-ICL on the WN18RR dataset is illustrated in Figure 6. Additionally, the hyper-parameters for models trained with the LoRA technique, specifically KG-LLAMA-7B, are outlined in Table 7.
It’s worth noting that for Table 2, Table 3, Table 4, Table 5 and Figure 4, we determine the most suitable layer for the task by experimenting with the validation set, and then report the results on the test set. For Figure 3, the results shown are for different layers of the language model on the test set. For the baseline results in Table 2 and Table 3, we give priority to using the results reported in the official baseline paper. For results not included in the paper, we replicate the experiments using the hyperparameters recommended by the official sources and conduct multiple attempts and optimizations on these hyperparameters as well. Each experiment is executed three times with random initializations, and the mean results are reported.
Parameter Value
optimizer adamw
learning rate 3e-4
epoch 3
weight decay 0
LoRA modules q_proj, k_proj, v_proj, o_proj
LoRA r 8
LoRA alpha 16
LoRA dropout 0.05
Table 7: Hyper-parameters of KG-LLAMA-7B.
Appendix B Additional Results and Analysis
B.1 Relation Prediction
Figure 7: Prompt templates to stimulate the LLM for relation prediction.
Method Procedure GPU Memory Time
KG-LLAMA
Training
(LLM:1 epoch)
14.68G
(LLM Parameters+LoRA+Gradient etc)
83h
(Forward+Backward)
Inference
(LLM:Generation)
12.94G
(LLM Parameters+LoRA)
2h50min
LLAMA-MLP
Training
(MLP:10 epoch)
0.078G
(MLP Parameters+Gradient etc)
33min
(Forward+Backward)
Inference
(LLM:Probing+MLP)
12.82G+0.078G
(LLM Parameters+MLP Parameters)
5h10min+29min+15s
Table 8: Efficiency Comparsion between KG-LLAMA and LLAMA-MLP on the WN11 dataset. The "Procedure" column represents several key steps of these methods.
Given an incomplete triple $(h,?,t)$ as query, relation prediction aims to predict the missing relation (denoted as ????). For relation prediction, we transform the binary framework of triple classification shown in Figure 1 into a multi-class framework. The categories for classification are the relationship types within each knowledge graph. Table 3 shows our experimental results on the YAGO3-10 (Dettmers et al., 2018) dataset. We use the following prompt template, $PT_{3}$ and $PT_{4}$, to create training samples for classification, as shown in Figure 7. $PT_{4}$ introduces the $D(e)$ to generate more precise and distinguishable hidden states, linking entities in the Yago with entities in the LLM, thereby improving the overall relation prediction performance.
It is worth noting that "ALL" represents 1,079,040 training samples. LLaMA-MLP and LLAMA-DES2 (GPT) only use 0.6% of the training set to train the probing classifier. It can be seen that LLAMA-DES2 (GPT) with frozen language model parameters achieves results very close to KG-LLAMA-7B, which is fine-tuned on the full dataset.
B.2 Case Study
Table 9 shows the prediction results of different models for two samples in the WN11 test set. For the first case, it can be seen that among the methods using frozen LLMs, only LLAMA-7B-ICL, LLAMA-MLP, and LLAMA-DES-H predict correctly. However, the LLAMA-7B-ICL method predicts "True" for almost all test samples. For the second case, it can be seen that only LLAMA-MLP, LLAMA-DES-H, and KG-LLAMA-7B predict correctly. However, KG-LLAMA-7B is fine-tuned on the full training set. The correct predictions by LLAMA-MLP and LLAMA-DES-H indicate that the LLM itself has the relevant knowledge to make accurate predictions, whereas methods like LLAMA-7B-ICL and ALPACA-7B do not fully utilize the knowledge stored in the language model or generate hallucinations during the response process.
QUERY MODEL PRED LABEL
HEAD: parsnip; REL: type of; TAIL: herb; LLAMA-7B N/A TRUE
LLAMA-7B-ICL TRUE
ALPACA-7B FALSE
ALPACA-7B-ICL FALSE
LLAMA-MLP TRUE
LLAMA-DES-H TRUE
KG-LLAMA-7B TRUE
HEAD: market synset; REL: domain topic; TAIL: brain stem; LLAMA-7B N/A FALSE
LLAMA-7B-ICL TRUE
ALPACA-7B TRUE
ALPACA-7B-ICL TRUE
LLAMA-MLP FALSE
LLAMA-DES-H FALSE
KG-LLAMA-7B FALSE
Table 9: Case study (triple classification) on two representative examples selected from the WN11 dataset. "N/A" indicates that the model’s output does not contain the predicted label information (i.e., yes or no).
B.3 Efficiency Analysis
We compared the performance of the KG-LLAMA method and LLAMA-MLP on the WN11 full dataset in terms of GPU memory and time consuming metrics using an Nvidia 3090 GPU. The results are shown in Table 8. For the computation of GPU memory usage, we use "torch.cuda.memory_allocated" to obtain the peak memory usage at each step. For the calculation of GPU memory reduction factor, we focus solely on the training phase, dividing the memory usage of KG-LLAMA by that of LLAMA-MLP. For the reduction factor of time consumption, we first sum the time consumed during both the training and inference phases for each method and then divide the results. For all methods, the batch size for all procedures is set to 1. The KG-LLAMA method employs LoRA for supervised fine-tuning, and its hyperparameters are listed in Table 7, with the epoch set to 1 in the efficiency experiment. For the training stage of KG-LLAMA, we employ 8-bit quantization technique, while for the inference stage, we configure not to use beam search. For the inference stage of LLAMA-MLP, we probe both the training data and the testing data. Thus, the sum of the three times in the table corresponds to the time required for probing the training data, probing the testing data, and the MLP forward pass, respectively.
It is worth noting that in practical use, our probing method only requires a small portion of the training set to complete the training of classifiers (such as MLP, Logistic Regression, etc.). Here, we used the full dataset for fairness in testing. We can clearly observe in the Table 8 that the proposed probing method, which uses frozen language models, consumes less GPU memory at all stages compared to KG-LLAMA, even though KG-LLAMA employs parameter-efficient fine-tuning techniques like LoRA to save memory. Additionally, since our method only requires a forward pass through the LLM, it consumes significantly less time compared to KG-LLAMA.
Method Pred Samples FB13 WN11 15K237N WN18RR UMLS
LLAMA-7B N-L 0 0.699 0.661 0.573 0.458 0.658
ALPACA-7B N-L 0 0.226 0.696 0.561 0.714 0.526
LLAMA-7B-ICL N-L 2 0.501 0.500 0.578 0.502 0.538
ALPACA-7B-ICL N-L 2 0.675 0.706 0.607 0.691 0.605
STRUCTURAL-AWARE-IT N-L ALL 0.907 0.959 0.752 0.927 0.881
ALPACA-DES N-L 0 0.611 - 0.590 0.548 0.525
LLAMA-MLP H-MLP 3k/10k 0.855 0.872 0.674 0.858 0.882
LLAMA-DES-H H-MLP 3k/10k 0.901 - 0.738 0.931 0.862
LLAMA-DES2 (GPT)* H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
LLAMA-DES-H H-MLP ALL 0.905 - 0.742 0.937 0.866
Table 10: The experiment results (accuracy) of triple classification. ALPACA-DES combines non-generated descriptions with $PT_{2}$ for Alpaca-7B.
B.4 Additional Baselines
From Table 10, it is noteworthy that ALPACA-DES fails to outperform ALPACA-7B, suggesting that mere incorporation of entity descriptions may not yield direct performance enhancements, considering the possibility of LLMs revealing the wrong responses and being influenced by hallucinations.
Meanwhile, in our study, we did not compare our method with KOPA because it integrates KG models (such as TransE/RotatE) with language models (like LLaMA) in a hybrid approach, thereby blending multiple approaches rather than solely utilizing the capabilities of LLMs (Zhang et al., 2023). In contrast, our current research exclusively focuses on the utilization of language models (like LLaMA), rendering KOPA an unsuitable baseline for our purposes. Nonetheless, we reproduced the Structural-aware-IT model proposed by Zhang et al. (2023), which incorporates in-context learning and instruction tuning. This model samples the neighborhoods of the head entity $h$ and tail entity $t$, including the textual descriptions of neighboring triples in the demonstration prompt, followed by fine-tuning on the entire training set. Like KOPA, this approach integrates structural information from the KG with LLM capabilities. Despite this, our method, which exclusively utilizes LLMs without integrating KG structural information, achieves performance comparable to the Structural-aware-IT model, which combines both KG structure and LLM capabilities.
B.5 Analysis of the Performance in the Intermediate Layers of LLM
The selection of the optimal layer index in LLMs serves as a hyperparameter, which we may select based on empirical evidence or adjust as needed when a validation set is available. Meanwhile, our experiments found a strong consistency for the same LLM across different knowledge graphs (for example, for LLAMA, the 15th layer often yields the best KGC performance), as shown in Figure 3, which reduces the complexity of utilization.
Additionally, the superior performance of the intermediate layers of language models, as compared to those closer to the output layer, may be attributed to the potential for each layer of the model to experience hallucinations. Moreover, each layer of a language model can potentially propagate hallucinations from previous layers, manifesting as deliberate falsehoods or the expression of misleading information (Zou et al., 2023). This phenomenon can lead to scenarios where models often know the correct answer even while generating incorrect outputs (Zou et al., 2023). Concurrently, our experimental results are consistent with the conclusions of Jiang et al. (2024), which found that the MLP layers in the later stages have an inhibitory effect on the extracted knowledge, particularly contributing significantly to erroneous output decoding. As a result, the middle layers in their experiments also tend to outperform the later layers.
B.6 Additional Analysis of the Data Efficiency
From Figure 8, we can observe the following: For the FB13 dataset, using 400 training samples, which is 0.06% of the training set, achieves 97.0% of the performance of using the entire training set; For the FB15K-237N dataset, using 0.57% of the training set achieves 99.6% of the performance of using the entire training set; For the WN18RR dataset, using 0.46% of the training set achieves 98.8% of the performance of using the entire training set; For the UMLS dataset, using 9.58% of the dataset achieves 90.6% of the performance of using the entire training set.
Figure 8: The LLAMA-DES-H prediction performance on KGC varies with the size of the training dataset. The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples, and the vertical axis represents the accuracy on the test set.
It can be observed that the LLAMA-DES-H method exhibits slightly lower data efficiency on the UMLS dataset compared to other methods. This phenomenon may be attributed to the fact that UMLS is a domain-specific knowledge graph, whereas the other datasets are encyclopedic knowledge graphs. The increased complexity of the decision boundary in the hidden states of the positive and negative example triplets generated by LLaMA necessitates a larger volume of data to adequately model the boundary.
B.7 Analysis of Methods for Entity Description Generation
From Table 2, it can be observed that the performance of LLAMA-DES2 (Tri) is significantly inferior to that of LLAMA-DES2 (GPT), and even LLAMA-MLP. This indicates that using subgraph triples as entity descriptions directly does not enhance the model’s classification performance; it may even impair its judgment. This phenomenon may be attributed to subgraph triples diverging significantly from the language model’s pre-training corpus, thus providing erroneous stimulus information. However, organizing triples into coherent semantic natural language text using methods in Appendix A.1 appears to mitigate this impact.
Figure 9: We visualize the hidden states obtained from stimulation in the FB13 test set in three-dimensional space using PCA. The high dimensional vectors are obtained using the LLAMA-DES2 (GPT) method, collecting the language model’s hidden states from the 16th layer through stimulation.
B.8 Analysis of Using Frozen LLM in KGC
It has been observed that the predictive performance achieved by the frozen language models can not achieve comparable performance with supervised fine-tuning the language models (Zhang et al., 2023; Yao et al., 2023). But the knowledge graphs used in the experiments are encyclopedic knowledge graphs, such as FB15K-237 (Toutanova and Chen, 2015), which are subsets of FB15k (Bordes et al., 2013) derived from Wikipedia. LLMs like LLaMA (Touvron et al., 2023), which are pretrained on datasets such as Wikipedia, CommonCrawl, etc., have stored this type of world knowledge (Davison et al., 2019). Liu et al. (2023) shows that prompting can effectively access the knowledge within LLMs without fine-tuning. Therefore, frozen LLMs have sufficient knowledge and capability to complete KGC tasks.