### v3

# Bridging LLMs and KGs without Fine-Tuning: Intermediate Probing Meets Subgraph-Aware Entity Descriptions

Abstract.
Traditional knowledge graph completion (KGC) methods rely solely on structural information, struggling with the inherent sparsity of knowledge graphs (KGs). By contrast, Large Language Models (LLMs) encapsulate extensive world knowledge and exhibit powerful context modeling capabilities, making them promising for mitigating the limitations of traditional methods. However, direct fine-tuning of LLMs for KGC, though effective, imposes substantial computational and memory overheads, while utilizing non-fine-tuned LLMs is efficient but yields suboptimal performance. In this work, we propose a novel framework that synergizes the strength of LLMs with robust knowledge representation to enable effective and efficient KGC. We extract the context-aware hidden states of knowledge triples from the intermediate layers of LLMs, thereby capturing rich semantic and relational nuances. These represenation are then utilized to train a data-effecient classifier tailored specifically for KGC tasks. To bridge the semanitic gaps between LLMs and KGs, we employ subgraph sampling on KGs to generate model-friendly entity descriptions, ensuring that the representations are comprehensive and well aligned with the KG structure. Extensive experiments on standard benchmarks validate the efficiency and effectiveness of our approach. We achieve a 47% relative improvement over previous methods based on non-fine-tuned LLMs and, to our knowledge, are the first to achieve classification performance comparable to fine-tuned LLMs while enhancing GPU memory efficiency by 188×188$\times$188 × and accelerating training and inference by 26.11×26.$1$.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Knowledge graph completion has become a crucial endeavor to improve the application of knowledge graphs (Chen et al., 2020), which aims to complete the knowledge graph and expand its scale by predicting the potential relationship between existing entities and discovering new relational facts (Shen et al., 2022b). Traditional knowledge graph embedding methods only utilize structure information from the knowledge graph (Bordes et al., 2013; Trouillon et al., 2016; Sun et al., 2019), which is vulnerable to the sparseness of KGs (Yao et al., 2019). By leveraging the powerful contextual modeling capabilities and extensive stored world knowledge of LLMs, we can effectively mitigate these challenges, and enhance the performance of KGC (Wei et al., 2024).
These approaches generally fall into two types: using non-fine-tuned LLMs with prompting techniques and fine-tuned LLMs specifically for KGC. The first type leverages LLMs with prompt techniques, such as in-context learning. These techniques enable models to perform KGC directly through natural language outputs (Yang et al., 2024; Wei et al., 2024; Yao et al., 2024), harnessing the pre-trained knowledge embedded within these models without additional training. While this approach offers computational efficiency, it may not fully capture the task-specific nuances required for optimal KGC performance. In contrast, the latter type involves fine-tuning the LLMs using supervised techniques to adapt models specifically for KGC tasks. This method involves fine-tuning LLMs, such as LLaMA, through the use of training samples (Yao et al., 2024), which can lead to improved performance by aligning the models more closely with the KGC objectives (Zhang et al., 2024; Yao et al., 2024). However, fine-tuning demands significant increases in time and resource expenditure. This dichotomy between computational efficiency and task-specific performance raises a fundamental question: How can we harness LLMs for effective and efficient KGC without incurring the high costs associated with fine-tuning?
Answering this question necessitates a deeper look at the limitations of LLMs compared to their fine-tuned counterparts. Non-fine-tuned LLMs, while cost-effective, risk generating misleading outputs in the form of hallucinations (Zhao et al., 2023) and underutilizing their latent knowledge (Wang et al., 2023). Moreover, the intricate relational structures and semantic subtleties embedded in KGs compound the challenge, since LLMs are not inherently designed to parse or disambiguate complex entity linkages, without explicitly fine-tuning on such tasks (Li et al., 2024).
In this paper, we propose a novel framework that reconciles these competing demands by fully exploiting the knowledge representation capabilities of LLMs without fine-tuning, as shown in Figure 1. Our method strategically extracts intermediate hidden states from LLMs to capture rich, context-aware representations of knowledge triples, which are then used to train a lightweight, data-efficient classifier specifically tailored for KGC. To bridge the gap between LLMs and KGs, we incorporate an innovative subgraph sampling technique to generate model-friendly entity descriptions, enabling the LLMs to better interpret and disambiguate relational patterns to further enhance knowledge representation. By focusing on these intermediate LLM representations, our framework mitigates hallucination risks while retaining the benefits of a non-fine-tuned paradigm, allowing for both high accuracy and reduced computational costs.
Extensive experiments demonstrate that our framework not only outperforms existing non-fine-tuned LLM-based methods by $47\%$ relative improvement but also achieves classification performance comparable to fine-tuned LLMs. Our approach enhances GPU memory efficiency by a factor of 188×188$\times$188 × and accelerates both training and inference processes by 26.11×26.$1$. Our contributions can be summarized as follows:
Figure 1. The overall architecture of our method for KGC involves categorizing triples based on the type of task, followed by their verbalization. This is combined with entity descriptions, which are generated by the subgraph entity description generator, to produce the stimulation prompt. Finally, the internal representation (hidden states) of the LLM obtained from the stimulation prompt is used to train a classifier, thereby completing the KGC. The only component that needs training in the entire architecture is the data-efficient classifier, making the method easy to train and cheap to deploy.
  • We propose, to the best of our knowledge, the first framework that directly harnesses the intermediate representation space of non-fine-tuned LLMs for KGC tasks. To bridge the semantic gap between the LLM and the KG, we design a subgraph-based entity description generator, ensuring that the model’s internal understanding of entities aligns effectively with KG structures.
  • We provide an in-depth analysis of why fine-tuned LLMs typically achieve superior KGC performance, identifying the core factors responsible for these gains. Building on these insights, we present novel optimization strategies for non-fine-tuned LLMs and are the first to enable them to achieve KGC performance comparable to their fine-tuned counterparts.
  • We perform extensive experiments across standard KGC datasets to validate the effectiveness and efficiency of our proposed method. These substantial improvements highlight the scalability and practical applicability of our method, making it a viable solution for large-scale KGC applications where computational resources and time are critical constraints.
2. Method
Our method is proposed in face of the limitations of non-fine-tuned LLMs on KGC. In Section 2.2, we present how to precisely extract context-aware representation of knowledge from intermediate layers of LLMs. In Section 2.3 we tackle complex relationships in KGs, by the generation of model-friendly entity descriptions through subgraph sampling.
2.1. Formulation and Notations
A knowledge graph, $G$, comprises a set of entities, $E$, and relations, $R$. It can be represented as a set of triples $G=\{(h,r,t)\}$, where $h\in E$ is the head entity, $t\in E$ is the tail entity, we collectively refer to the head entity and tail entity as $e$, and $r\in R$ represents the relation between them. $D$ is the set of descriptions for each entity and relation. We denote $D(e)$, $D(r)$ as the textual description of each entity and each relation.
For triple classification, the task is to assess the plausibility of a given knowledge triple, determining whether the triple is true or not. As for relation prediction and entity prediction, they aim to predict the missing relation and entity in a triple, respectively.
2.2. Extracting Knowledge Representations
Knowledge graphs are composed of numerous triples that encapsulate complex relational and semantic information. For non-fine-tuned LLMs, interpreting these intricate structures poses significant challenges, often leading to the retrieval of irrelevant information and the generation of incorrect or misleading responses.
To enhance the performance of LLMs in KGC, our method leverages the internal judgment mechanisms of LLMs, as shown in (Zou et al., 2023). Specifically, we strategically harness the intermediate layers of LLMs to extract rich, context-aware knowledge representations that are pivotal for accurate KGC.
First, we construct specialized prompt templates designed to stimulate the LLM to generate the necessary classification information within its intermediate layers. For triple classification, where the objective is to determine the plautibility of triples in the KG, we employ a prompt template, $PT_{1}$, that creates positive and negative sample pairs for training, as illustrated in Figure 2. Notably, the prompt avoids explicitly describing the positivity or negativity of triples.
Thereafter, we use existing triples in the knowledge graph to construct positive samples $S^{+}$, and employ the negative sampling technique (Lv et al., 2022) to obtain negative samples $S^{-}$. This results in a dataset obtaining $N$ training samples. Each training sample is input into the LLM to obtain representations from its intermediate layers.
Specifically, we capture plausibility information and extract hidden states for positive samples $s^{+}\in S^{+}$ as follows:
(1) $Rep_{j}^{i}\left(s^{+}\right)=\{M_{1:i}\left(s^{+}\right)\left[j\right],0\textless i\textless L,0\textless j\textless N\},$
where $M$ represents the language model, $M_{1:i}$ represents the first $i$ layers of the model, $[j]$ refers to the hidden state corresponding to the $j$th token, $L$ and $N$ represent the number of layers in the model and the number of tokens resulting from tokenizing $s^{+}$. Consistent with findings in (Zou et al., 2023), we utilize the representation of the last token in each input text, as it is more conducive to information extraction due to the autoregressive nature of causal language models. Negative samples are processed similarly to ensure consistency in representation extraction.
The extracted representations $Rep_{j}^{i}(s^{+})$ and $Rep_{j}^{i}(s^{-})$ serve as input features for training a lightweight classifier to distinguish between plausible and implausible triples by learning a decision boundary that effectively separates the hidden states of positive samples from those of negative samples. Although the classifiers employed—such as multilayer perceptrons (MLPs) and logistic regression models—are inherently simple, their effectiveness is significantly amplified by the quality and richness of the intermediate layer representations, thereby enhancing the ability to make accurate predictions without the computational overhead associated with fine-tuning the entire LLM.
For relation prediction, we extend our binary classification framework to a multi-class setting, where the classifier predicts the specific relationship types within the KG. Utilizing the $PT_{3}$ shown in Figure 2, we stimulate the LLM to generate hidden states that capture the characteristics of different relation types. This approach ensures that the classifier can effectively learn to differentiate between multiple relationship categories from the model’s internal processing, enhancing the accuracy of KGC.
Figure 2. Prompt templates to stimulate the LLM.
2.3. Subgraph Entity Description Generator
The semantics of entities in a KG can be complex and ambiguous, posing significant challenges for LLMs to interpret accurately based solely on entity names. To overcome these challenges and enhance the semantic comprehension of entities by LLMs without resorting to computationally intensive fine-tuning, we introduce the subgraph entity description generator, which enriches entity representations by providing additional contextual descriptions derived from their local subgraph structures. Specifically, we extend the prompt template constructed in Section 2.2. We focus on aligning the language model’s internal entity understanding with those in the knowledge graph. Therefore, we construct new prompts $PT_{2}$ and $PT_{4}$, as shown in Figure 2. By introducing the entity description $D(e)$, LLMs are able to generate more precise and distinguishable hidden states, thereby improving the overall classification performance.
Most KGs contain entity descriptions that can be directly used as $D(e)$. However, to address incomplete entity descriptions within certain KG datasets, we propose a method to generate $D(e)$. Unlike the approach in (Yang et al., 2024), which generates the entity description using only a single triple, our method constructs a more precise $D(e)$ by incorporating subgraph information in two distinct ways. For the first approach, we concatenate all the triples within the one-hop subgraph around the entity $e$, thereby forming $D(e)$ (Zhang et al., 2024). However, providing information in the form of explicit rules may be less effective for LLM to retrieve and leverage knowledge (Bian et al., 2024), as the structure of these subgraph triples may significantly differ from the text seen during the pre-training phase. Therefore, our second approach leverages a language model to transform these triples into a narrative format that is more comprehensible to the LLM. By rephrasing the triples from the one-hop subgraph around the entity $e$ into natural language text, we produce a more refined and model-friendly version of $D(e)$ that is more effective for LLMs to retrieve and leverage knowledge. Specifically:
(2) $D(e)=\operatorname{Rephrase}(\{\operatorname{Verbalize}(\{h_{i}\}\{r_{i}\}\{t_{i}\}),$
$(h_{i},r_{i},t_{i})\in\operatorname{Subgraph}(e)\}),$
where $\operatorname{Subgraph(\cdot)}$ is a function that automatically retrieves a specified number of triples in the one-hop subgraph where the entity resides; $\operatorname{Verbalize(\cdot)}$ is a function that transform the triple into its textual string, in this work, we use the linear verbalization (Baek et al., 2023); For the first approach, $\operatorname{Rephrase(\cdot)}$ is a string concatenation function that joins each verbalized triple statement with a separator. For the second approach, we employ a LLM with the in-context learning technique, then ask it to generate model-friendly descriptions of entities based on the verbalized triples in their subgraph.
By leveraging the subgraph entity description generator, we align the LLM’s internal entity understanding with those in the KG, akin to entity linking process. This alignment not only mitigates ambiguities but also enhances LLM’s semantic comprehension, ultimately boosting KGC performance.
Method Pred Samples FB13 WN11 15K237N WN18RR UMLS
TransE S-F ALL 0.815 0.759 0.697 0.884 0.845
DistMult S-F ALL 0.862 0.871 0.587 0.851 0.864
ComplEx S-F ALL 0.857 0.862 0.657 0.841 0.871
RotatE S-F ALL 0.819 0.847 0.685 0.882 0.877
KG-BERT L-MLP ALL 0.899 0.935 0.560 0.916 0.773
KGT5 N-L ALL 0.663 0.728 - - -
LLAMA-7B N-L 0 0.699 0.661 0.573 0.458 0.658
LLAMA-7B-ICL N-L 2 0.501 0.500 0.578 0.502 0.538
KG-LLAMA-7B N-L ALL 0.892 0.955 0.748 0.921 0.858
KG-LLAMA-SAMPLED N-L 3k/10k 0.727 0.682 0.614 0.574 0.834
LLAMA-MLP H-MLP 3k/10k 0.855 0.872 0.674 0.858 0.882
LLAMA-DES-H H-MLP 3k/10k 0.901 - 0.738 0.934 0.862
LLAMA-DES2 (Tri) H-MLP 3k/10k 0.842 0.755 0.671 0.779 0.782
LLAMA-DES2 (GPT) H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
LLAMA-DES-H H-MLP ALL 0.920 - 0.753 0.943 0.863
Table 1. The main experiment results (accuracy) of triple classification. In the ‘Pred’ column, $S\mbox{-}F$ uses score function, $L\mbox{-}MLP$ uses MLP on last layer hidden states, $N\mbox{-}L$ uses natural language, and $H\mbox{-}MLP$ uses MLP on intermediate layer hidden states to predict. ‘Samples’ indicates training sample counts, with $3k/10k$ indicating $10k$ samples for the UMLS dataset and $3k$ for others. LLAMA-MLP uses LLaMA-7B with MLP for hidden state classification and $PT_{1}$ for stimulus generation; LLAMA-DES-H combines non-generated descriptions with $PT_{2}$ for stimulus generation; LLAMA-DES2 (Tri)/(GPT) uses first/second generated descriptions with $PT_{2}$ for stimulus generation. Top three results per dataset are bolded.
3. Experiments
In the experiments, our objective is to answer the following research questions (RQs):
  • RQ1: Can LLMs with our method effectively achieve performance on par with their fine-tuned counterparts?
  • RQ2: How does our method perform across different language models and classification models?
  • RQ3: What are the data efficiency and computation efficiency of our proposed method for KGC?
3.1. Datasets
Dataset $|\mathbf{E}|$ $|\mathbf{R}|$ Train # Valid # Test
FB13 75,043 13 316,232 5,908 23,733
WN11 38,696 11 112,581 2,609 10,544
FB15K-237N 13,104 93 87,282 14,082 16,452
WN18RR 40,943 11 86,835 6,068 6,268
UMLS 135 46 5,216 1,304 1,322
YAGO3-10 123,182 37 1,079,040 5,000 5,000
Table 2. Statistical information of datasets.
In our experiments, we employ five widely-used KG datasets for triple classification: FB13 (Socher et al., 2013), WN11 (Socher et al., 2013), FB15K-237N (Lv et al., 2022), WN18RR (Dettmers et al., 2018), and UMLS (Yao et al., 2019). For relation prediction, we employ the YAGO3-10 dataset (Dettmers et al., 2018), and for entity prediction, we use WN18RR for evaluation.
For FB13, we utilize Wikipedia to construct entity descriptions. For FB15K-237N, the entity descriptions provided by (Xie et al., 2016) are used. Synset definitions from (Yao et al., 2019) served as entity descriptions for WN18RR. In the case of UMLS, entity descriptions provided by the dataset are used. In the following, we refer to these entity descriptions as non-generated entity descriptions.
3.2. Experimental Settings
To generate entity descriptions, for the first approach in Section 2.3, we concatenate ten randomly selected triples from the one-hop subgraph surrounding each entity to construct the descriptions. For the second approach, we employ GPT-3.5-turbo as the LLM. In the following text, we refer to these entity descriptions as generated entity descriptions (Tri)/(GPT).
We compare our method with four structure information-based methods: TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), ComplEx (Trouillon et al., 2016) and RotatE (Sun et al., 2019), as well as seven additional information-based methods: KG-LLAMA (Yao et al., 2024), KG-BERT (Yao et al., 2019), KG-T5 (Saxena et al., 2022), LLaMA-7b, Alpaca-7B, LLaMA-7B-ICL and Alpaca-7B-ICL (Zhang et al., 2024). KG-LLAMA, KG-BERT and KGT5 fine-tune the language models using all the training samples. LLaMA-7B and Alpaca-7B rely on non-fine-tuned LLMs to directly perform predictions. LLaMA-7B-ICL and Alpaca-7B-ICL employ LLMs with in-context learning to make predictions (Zhang et al., 2024). We give priority to using the results reported in the official baseline paper. Please refer to Section 4.1 for detailed introductions of these baseline methods.
For the negative sampling employed on the datasets, we utilize the same negative sampling technique as (Yao et al., 2019). For the intermediate layer selection, we identify the most suitable layer for the task by experimenting with the validation set. Each experiment is executed three times with random initializations, and the mean results are reported.
3.3. Main Results (For RQ1)
From Table 1, it can be observed that LLAMA-DES-H and LLAMA-DES2 (GPT), when using a partial training set, exhibit comparable performance to KG-LLAMA-7B, the latter being fine-tuned with all training triples. Moreover, when trained on subsets of equivalent size, LLAMA-DES-H and LLAMA-DES2 (GPT) markedly surpass KG-LLAMA-SAMPLED, the existing state-of-the-art approach that relies exclusively on LLaMA for KGC.
The experimental results presented in Table 1 also validate the effectiveness of the approaches proposed in Section 2.2 and Section 2.3. Specifically, LLAMA-DES-H and LLAMA-DES2 (GPT) outperform LLAMA-MLP, demonstrating the effectiveness of using subgraph entity description generator to improve the LLM’s understanding of entities in the KG, as proposed in Section 2.3. Furthermore, LLAMA-MLP demonstrates superior performance compared to LLAMA-7B and LLAMA-7B-ICL, affirming the effectiveness of stimulation method proposed in Section 2.2. Thus, the experimental results substantiate the effectiveness of our proposed methodology. Additionally, the non-fine-tuned LLM can also achieve similar KGC effects as that of fine-tuning the LLM.
Notice that the performance on the UMLS dataset is special, where the use of generated descriptions hinders the model’s classification accuracy, and the use of non-generated entity descriptions is also worse than not using any entity descriptions. It is postulated that the underlying reason may lie in the highly specialized and domain-specific nature of entity names in the UMLS dataset, which are often precise medical terms with limited room for multiple interpretations. The entities in the knowledge graph have well-aligned with the internal entities of the LLM. Meanwhile, using entity descriptions may be affected by noisy description text, and descriptions generated by GPT-3.5-turbo may be influenced by hallucination, thereby affecting the classification effectiveness of the induced hidden states. Meanwhile, our method shows slightly lower performance on the WN11 and FB15K237N datasets compared to KG-LLAMA, which fine-tuned on the full training set. This discrepancy may stem from the fact that the pretrained LLM did not acquire knowledge of certain triples in the KG during its pretraining phase, resulting in incorrect predictions even when probing is conducted.
In addition to the triple classification, we also validate the effectiveness of the proposed probing method on the relation prediction task as shown in Table 3. The experimental results similarly demonstrate that the probing method can achieve the nearly same performance as full training set fine-tuning, even when using only 0.6% of the training data. For entity prediction, we obtain the plausibility confidence scores of the triples through the classifier and rank them accordingly. Table 4 also displays similar results, showing that after using model-friendly descriptions, the LLAMA-DES2 (GPT) method improves the performance of LLAMA-MLP-7B and surpasses its fine-tuned counterparts.
Figure 3. Differences in KGC performance across different layers of the LLAMA-DES2 (GPT) [left] and Mistral-DES2 (GPT) [right]. The horizontal axis represents the position of the intermediate layers in LLaMA or Mistral, and the vertical axis represents the dataset.
Method Pred Samples YAGO3-10
KG-BERT L-MLP ALL 0.6816
KGT5 N-L ALL 0.5714
ChatGLM-6B N-L 0 0.0658
KG-ChatGLM-6B N-L ALL 0.5662
LLaMA-7B N-L 0 0.0348
LLaMA-13B N-L 0 0.0040
KG-LLaMA-7B N-L ALL 0.7028
LLAMA-MLP-7B H-MLP 6996 0.5968
LLAMA-DES2 (GPT) H-MLP 6996 0.6824
LLAMA-DES2 (GPT) H-MLP ALL 0.7015
Table 3. The experiment results (Hits@1) of relation prediction.
Method Pred Samples WN18RR
KG-BERT L-MLP ALL 0.1102
StAR L-MLP ALL 0.2430
KGT5 N-L ALL 0.1011
KG-ChatGLM-6B N-L ALL 0.1613
LLaMA-7B N-L 0 0.0849
LLaMA-13B N-L 0 0.0991
KG-LLaMA-7B N-L ALL 0.2415
LLAMA-MLP-7B H-MLP 10k 0.2108
LLAMA-DES2 (GPT) H-MLP 10k 0.2433
LLAMA-DES2 (GPT) H-MLP ALL 0.2495
Table 4. The experiment results (Hits@1) of entity prediction.
3.4. Ablation Study (For RQ2)
Through the analysis depicted in the left portion of Figure 3, it is observed that with an increase in the index of the intermediate layer, the efficacy of hidden states for triple classification demonstrates an initial incline followed by a subsequent decline. Intermediate layers closer to natural language text output generally exhibit better prediction performance compared to the lowest layers. This observation corroborates the notion posited by (Geva et al., 2020) that language models encode knowledge differentially across their Feed-Forward layers, with higher layers integrating information from lower layers. During pretraining, the lower layers may not have stored certain knowledge, and thus are unable to generate hidden states conducive to accurate triple classification. Furthermore, the prediction performance of the layers closer to the natural language text output is generally worse than those in the middle layers. This phenomenon may be attributed to the hallucination issue inherent in LLaMA, wherein higher layers could potentially either omit or alter triple information, leading to a blurred distinction between positive and negative triple hidden states induced by stimuli (Jiang et al., 2024; Zou et al., 2023). Akin observations are discernible on the right side of Figure 3 for the Mistral model, which has a different backbone from LLaMA. This experimental results are consistent with the conclusions of (Jiang et al., 2024), which found that the MLP layers in the later stages have an inhibitory effect on the extracted knowledge, particularly contributing significantly to erroneous output decoding. As a result, the middle layers in their experiments also tend to outperform the later layers. Meanwhile, as observed in Table 6, our method demonstrates strong versatility across different backbone models, achieving performance comparable to full dataset fine-tuning with only a small sample size.
Regarding classification models, a comparative analysis encompassing MLP, SVM, and Logistic Regression, as delineated in Table 5, is conducted. The result shows that MLP outperforms all three datasets compared to the other two classification models, while Logistic Regression also exhibits results close to MLP. This underscores the efficacy of our stimulus methodology, whereby the model inherently encodes discriminative features for KGC tasks, making hidden states easily distinguishable in high-dimensional space and allowing straightforward classification boundaries to emerge without the need for complex, nonlinear decision functions.
Method FB13 WN11 FB15K-237N
LLAMA-SVM 0.821 0.866 0.594
LLAMA-LR 0.839 0.860 0.662
LLAMA-MLP 0.855 0.872 0.674
Table 5. The variation in KGC prediction performance across different datasets using SVM, Logistic Regression, MLP as classification models. All these methods use LLaMA-7B as the base model, and generate stimuli using $PT_{1}$.
Model Method FB13 WN11
LLaMA KG-LLAMA-7B 0.892 0.955
LLAMA-MLP 0.855 0.872
LLAMA-DES-H 0.901 -
LLAMA-DES2 (GPT) 0.912 0.917
Mistral KG-MISTRAL-7B 0.891 0.962
MISTRAL-MLP 0.846 0.901
MISTRAL-DES-H 0.880 -
MISTRAL-DES2 (GPT) 0.893 0.925
Gemma KG-GEMMA-7B 0.897 0.944
GEMMA-MLP 0.806 0.893
GEMMA-DES-H 0.891 -
GEMMA-DES2 (GPT) 0.867 0.917
Table 6. The variation in KGC prediction performance across different datasets using LLaMA, Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024) as base models (which have different backbones). Please refer to Table 1 for the number of training samples used by different methods.
Method Procedure GPU Memory Time
KG-LLAMA
Training
(LLM:1 epoch)
14.68G
(LLM Parameters+LoRA+Gradient etc)
83h
(Forward+Backward)
Inference
(LLM:Generation)
12.94G
(LLM Parameters+LoRA)
2h50min
LLAMA-MLP
Training
(MLP:10 epoch)
0.078G
(MLP Parameters+Gradient etc)
33min
(Forward+Backward)
Inference
(LLM:Probing+MLP)
12.82G+0.018G
(LLM Parameters+MLP Parameters)
2h44min+15s
(Probing+MLP Forward)
Table 7. Efficiency Comparsion between KG-LLAMA-7B and LLAMA-MLP on the WN11 full dataset. The ‘Procedure’ column represents several key steps of these methods.
3.5. Efficiency Study (For RQ3)
As illustrated in Figure 4, our method outperforms both KG-LLAMA and KG-BERT on the FB13 when trained with only 100 samples, as opposed to the 3000 samples used by the latter methods. For the UMLS, the performance of KG-LLAMA and KG-BERT with 3000 training samples is comparable to that of LLAMA-DES-H trained with only 500 samples. Additionally, combining the results from Table 1, for the FB13 dataset, our method attains 97.2% of the performance achievable with the full training set, utilizing only 500 samples, which constitutes a mere 0.079% of the total training samples. In contrast, KG-LLAMA and KG-BERT achieve only 78.7% and 63.1% of the maximum performance, respectively. For the UMLS, with 9.58% of the training set, our method reaches 90.6% of the full training set performance. Our experimental results valid that the language model is capable of generating effective hidden states through our method, demonstrating its potential for data-efficient scenarios within the context of KGC tasks.
In addition to data efficiency, we also evaluate the computational efficiency of our method. Table 7 shows that our method achieves significantly reduced resource usage compared to parameter-efficient fine-tuning. For the computation of GPU memory usage, we use ”torch.cuda.memory_allocated” to obtain the peak memory usage at each step. The KG-LLAMA method employs LoRA and 8-bit quantization technique for supervised fine-tuning. Compared to fine-tuned 7B LLMs, we boost the GPU memory efficiency to 188×188$\times$188 × during model training and speed up 26.11×26.$1$ overall. Notably, the training overhead of our method lies solely in the classifier, which is independent of the LLM size. Therefore, when scaling LLMs to larger sizes, such as 70B, the substantial overhead associated with backpropagation in fine-tuning becomes increasingly burdensome. In these scenarios, the advantages of our method in terms of GPU memory usage and time consumption become even more pronounced.
Figure 4. The prediction performance of KG-LLAMA, KG-BERT, and LLAMA-DES-H on triple classification varies with the size of the training dataset. The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples.
Figure 5. Generalization across Datasets, LLAMA-DES-H. The horizontal axis represents the training set, and the vertical axis represents the test set.
Figure 6. Performance of LLAMA-DES2 (GPT) versus the number of sampled triplets.
4. Related Work
4.1. Knowledge Graph Completion
Knowledge graphs are typically characterized by their inherent incompleteness (Hogan et al., 2021). The prevailing methodologies in KGC can be categorized into two main classes: structure information-based methods and additional information-based methods (Shen et al., 2022b). Structure information-based methods leverage the inherent structural data present within knowledge graphs. For example, RESCAL (Nickel et al., 2011), DistMult (Yang et al., 2014), COMPGCN (Vashishth et al., 2019), and ConvE (Dettmers et al., 2018) primarily use scoring functions to evaluate the semantic similarities of entity or relation embeddings within a latent embedding space to infer missing information. Additional information-based methods incorporate a variety of supplementary data to enrich the knowledge graphs. The prominent examples are KG-BERT (Yao et al., 2019) and LASS (Shen et al., 2022a), which represent entities, relations, and triples as textual sequences, turning KGC into a sequence classification problem. KG-S2S (Chen et al., 2022) and KGT5 (Saxena et al., 2022) employ encoder-decoder architecture language models to complete KGC in a generative manner. And (Li et al., 2024) enhances KGC with LLM by generating entity descriptions and designing loss functions for fine-tuning. More recently, KG-LLAMA (Yao et al., 2024) is the first to adopt a decoder-only LLM architecture for KGC by formulating the task as question answering through an instruction-tuning-based prompt, followed by parameter-efficient fine-tuning on LLaMA. Notably, KG-LLAMA leverages only LLaMA’s own capabilities without incorporating additional hybrid approaches, which aligns with the core principle of our proposed framework.
4.2. Large Language Models
Several studies have explored the mechanisms through which Large Language Models process, interpret, and generate factual content. As highlighted by (Saier et al., 2023), these models function as extensive repositories of knowledge, encapsulating a vast array of information about the world within their parameters. This insight into LLMs underscores their potential as powerful tools for a wide range of applications, including knowledge graph completion, by leveraging the rich informational content stored in their intricate network structures. (Ramrakhiyani et al., 2023) have conducted evaluations across multiple Large Language Models to assess their knowledge of world geography using probing techniques. Similarly, (Jin et al., 2024) used probing to measure the abilities of different LLM layers to learn various levels of concepts. In addition, (Liu et al., 2024) investigated whether a universal truthfulness hyperplane exists within LLMs by employing probing methods.
5. Conclusion
This paper proposes a novel framework that leverages the latent representation space of LLMs to improve KGC. By integrating subgraph-based entity descriptions, our approach bridges the semantic gap between LLMs and KGs, enabling LLMs to achieve performance on par with their fine-tuned counterparts. Simultaneously, our method enhances data efficiency and saves a significant amount of GPU memory and time compared to parameter-efficient fine-tuning. This work explores the potential of LLMs in KGC tasks with both efficacy and efficiency. Compared to fine-tuning, we hit the core of KGC problem in a better way.